# Interaction of the Primitive Equations with Sea Ice Dynamics

### v3

@@ 4.1. The hydrostatic Helmholtz projection and the hydrostatic Stokes operator
(4.1) $\partial_{t}v-\Delta v+({u}\cdot\nabla){v}+\nabla_{\mathrm{H}}\pi&=f,&&\text{ on }{\Omega\times[0,T]},\\ \partial_{z}\pi&=0,&&\text{ on }{\Omega\times[0,T]},\\ \mathrm{div}\,u&=0,&&\text{ on }{\Omega\times[0,T]},\\ v(0)&=u_{0},&&\text{ on }{\Omega}.\end{aligned}\right.$
@@
We now introduce the associated hydrostatic Helmholtz or Leray projection $\mathcal{P}\colon\mathrm{L}^{q}(\Omega)^{2}\to{\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)$. It is a bounded, linear projection which annihilates the pressure term, i. e., $\mathcal{P}(\nabla_{\mathrm{H}}\pi)=0$, and the space $\mathrm{L}^{q}(\Omega)$ can be decomposed into $\mathrm{L}^{q}(\Omega)^{2}={\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)\oplus\{\nabla_{\mathrm{H}}\pi\colon\pi\in\widehat{\mathrm{W}}^{1,q}(G)\}$ along the hydrostatic Helmholtz projection $\mathcal{P}$. Furthermore, the Helmholtz projection commutes with the time derivative, i. e., $\partial_{t}\mathcal{P}=\mathcal{P}\partial_{t}$. It is related to the two-dimensional classical Helmholtz projection $\mathcal{P}_{\mathrm{H}}$, see e. g. [27, Section 2] for further details on the latter one, by
@@ 4.2. Hibler’s sea ice operator
Proceeding as in [9, Section 3], we introduce the map $\mathbb{S}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}^{2\times 2}$ with
@@
from [9, Section 3], where $i=1,2$ and $\mathrm{D}_{m}=-\mathrm{i}\partial_{m}$. The principal part of $\mathcal{A}^{\mathrm{H}}$ has coefficients given by
@@
$[\mathcal{A}^{\mathrm{H}}(v_{0}){v}_{\mathrm{ice}}]_{i}=\sum_{j,k,l=1}^{2}a_{ij}^{kl}(\varepsilon_{0},h_{0},a_{0})\mathrm{D}_{k}\mathrm{D}_{l}v_{\mathrm{ice},j}+\frac{1}{2{\rho}_{\mathrm{ice}}h_{0}\triangle_{\delta}(\varepsilon_{0})}\sum_{j=1}^{2}(\partial_{j}P(h_{0},a_{0}))(\mathbb{S}\varepsilon({v}_{\mathrm{ice}}))_{ij}.$
The $\mathrm{L}^{q}$-realization $\mathrm{A}^{\mathrm{H}}(v_{0})$ of $\mathcal{A}^{\mathrm{H}}(v_{0})$ on $G=(0,1)\times(0,1)$ is then defined by $[\mathrm{A}^{\mathrm{H}}(v_{0})]{v}_{\mathrm{ice}}:=[\mathcal{A}^{\mathrm{H}}(v_{0})]{v}_{\mathrm{ice}}$ for ${v}_{\mathrm{ice}}\in D(\mathrm{A}^{\mathrm{H}}(v_{0}))=\mathrm{H}_{\mathrm{per}}^{{2},q}(G)^{2}$.
Comparing this definition of Hibler’s operator to the one presented in [9, Section 3], we observe that the sign is changed and $\nicefrac{{1}}{{{\rho}_{\mathrm{ice}}h}}$ is included in the definition for the sake of consistency with the hydrostatic Stokes operator and to ease notation. We also introduce the lower-order terms
@@ 4.3. A new formulation of the coupled system
$\mathrm{A}^{\mathrm{atm}}{v}_{\mathrm{atm}}&={\mathcal{P}}_{\mathrm{atm}}\Delta{v}_{\mathrm{atm}},\\ D(\mathrm{A}^{\mathrm{atm}})&=\left\{{v}_{\mathrm{atm}}\in\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}}):\partial_{\nu_{\Gamma_{u}}}{v}_{\mathrm{atm}}=0\text{ on }{\Gamma_{u}}\text{ and }\partial_{\nu_{\Gamma_{i}}}{v}_{\mathrm{atm}}=0\text{ on }{\Gamma_{i}}\right\}\end{aligned}\right.$
@@
$\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}{v}_{\mathrm{ocn}}={\mathcal{P}}_{\mathrm{ocn}}\Delta{v}_{\mathrm{ocn}},\enspace D(\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}})=\{{v}_{\mathrm{ocn}}\in{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}}):{\mathcal{P}}_{\mathrm{ocn}}\Delta{v}_{\mathrm{ocn}}\in{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\},$
@@
(4.3) $\mathrm{A}^{\mathrm{ocn}}_{0}{v}_{\mathrm{ocn}}&={\mathcal{P}}_{\mathrm{ocn}}\Delta{v}_{\mathrm{ocn}},\\ D(\mathrm{A}^{\mathrm{ocn}}_{0})&=\left\{{v}_{\mathrm{ocn}}\in\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}}):\mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}=0\text{ on }{\Gamma_{b}}\text{ and }\mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}=0\text{ on }{\Gamma_{o}}\right\}\end{aligned}\right.$
@@
(4.4) $\mathcal{A}(v_{\mathrm{ice},0},h_{0},a_{0})$ $:=\begin{pmatrix}\mathrm{A}^{\mathrm{atm}}&0&0&0&0\\ 0&\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}&0&0&0\\ 0&-C_{\mathrm{o,i}}(h_{0})\partial_{\nu_{\Gamma_{o}}}&\mathrm{A}^{\mathrm{H}}(v_{\mathrm{ice},0},h_{0},a_{0})&\mathrm{B}_{h}(h_{0},a_{0})&\mathrm{B}_{a}(h_{0},a_{0})\\ 0&0&0&d_{\mathrm{h}}\Delta_{\mathrm{H}}&0\\ 0&0&0&0&d_{\mathrm{a}}\Delta_{\mathrm{H}}\end{pmatrix},$
$\mathcal{X}_{1}:=D(\mathcal{A})$ :=$:=\{({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)\in D(\mathrm{A}^{\mathrm{atm}})\times D(\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}})\times D(\mathrm{A}^{\mathrm{H}})\times D(\Delta_{\mathrm{H}})\times D(\Delta_{\mathrm{H}})\colon$
@@
[…]
@@ 5.1. Bounded $\mathcal{H}^{\infty}$-calculus of Hibler’s operator
The embedding below is classical, see e. g. [49, Theorem 4.6.1]. Let $p,q\in(1,\infty)$ with $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$. Then
@@
Making use of (5.1) and proceeding as in [9, Section 4], we obtain the ellipticity of $-\mathcal{A}^{\mathrm{H}}(v_{0})$. For the notion of strong ellipticity and parameter ellipticity, we refer e. g. to [13, Definition 5.1].
[…]
Let $p,q\in(1,\infty)$ such that $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$, and let $u_{0}=(v_{\mathrm{ice},0},h_{0},a_{0})\in V^{\mathrm{ice}}$. Then the principal part of the negative Hibler operator $-\mathcal{A}^{\mathrm{H}}(v_{0})$ is strongly elliptic and also parameter-elliptic of angle $\varphi_{-\mathcal{A}^{\mathrm{H}}(v_{0})}=0$.
We will now derive the bounded $\mathcal{H}^{\infty}$-calculus for the linearized $\mathrm{L}^{q}$-realization $-\mathrm{A}^{\mathrm{H}}(v_{0})$.
[…]
Let $p,q,r,s\in(1,\infty)$ such that $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$, let $v_{0}=(v_{\mathrm{ice},0},h_{0},a_{0})\in V^{\mathrm{ice}}$, and let $\mathrm{A}^{\mathrm{H}}(v_{0})$ denote the $\mathrm{L}^{q}$-realization of the linearized Hibler operator. Then there is $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, $-\mathrm{A}^{\mathrm{H}}(v_{0})+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathrm{L}^{r}(G)^{2}$ with angle $\Phi_{-\mathrm{A}^{\mathrm{H}}(v_{0})+\omega}^{\infty}<\nicefrac{{\pi}}{{2}}$.
@@
Identifying functions on the square $G=(0,1)\times(0,1)$ endowed with periodic boundary conditions with functions on the torus $\mathbb{T}$, exploiting Subsection 5.1 for the ellipticity properties of the negative Hibler operator $-\mathcal{A}^{\mathrm{H}}(v_{0})$ and using (5.1) in conjunction with the fact that the coefficients of $\mathcal{A}^{\mathrm{H}}(v_{0})$ depend smoothly on $\nabla_{\mathrm{H}}v_{\mathrm{ice},0}$, $\nabla_{\mathrm{H}}h_{0}$, $\nabla_{\mathrm{H}}a_{0}$, $h_{0}$ and $a_{0}$, we conclude from [15, Theorem 7.1] that there is $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, the operator $-\mathrm{A}^{\mathrm{H}}(v_{0})+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathrm{L}^{r}(\mathbb{T})^{2}=\mathrm{L}^{r}(G)^{2}$ with angle $\Phi_{-\mathrm{A}^{\mathrm{H}}(v_{0})+\omega}^{\infty}<\nicefrac{{\pi}}{{2}}$. ∎
The following result is a direct consequence of Subsection 5.1, because the latter proposition implies that $-\mathrm{A}^{\mathrm{H}}(v_{0})+\omega$ has bounded imaginary powers on $\mathrm{L}^{q}(G)^{2}$, see also the relation (8.7).
[…]
@@ 5.2. Estimates of the nonlinearities
(5.2) $\mathrm{S}_{\mathrm{h}}=f_{1}(\nicefrac{{h}}{{a}})a+(1-a)f_{1}(0)\quad\text{and}\quad\mathrm{S}_{\mathrm{a}}=\begin{cases*}\frac{f_{1}(0)}{\kappa_{1}}(1-a),&if $f_{1}(0)>0$,\\ 0,\quad&if $f_{1}(0)<0$,\end{cases*}\quad+\quad\begin{cases*}0,&if $\mathrm{S}_{\mathrm{h}}>0$,\\ \frac{a}{2h}\mathrm{S}_{\mathrm{h}},&if $\mathrm{S}_{\mathrm{h}}<0$,\end{cases*}$
@@
(5.3) $\mathcal{Y}_{0}&={\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})\times{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{4},\\ \mathcal{Y}_{1}&=\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})\times\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{H}_{\mathrm{per}}^{{2},q}(G)^{4},\\ \mathcal{Y}_{\theta}&=\mathrm{B}^{{2\theta}}_{{q}{p},\mathrm{per}}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})\times\mathrm{B}^{{2\theta}}_{{q}{p},\mathrm{per}}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{B}^{{2\theta}}_{{q}{p},\mathrm{per}}(G)^{4},\\ \mathcal{Y}_{\gamma}&=\mathcal{Y}_{\theta}^{\mathrm{atm}}\times\mathcal{Y}_{\theta}^{\mathrm{ocn}}\times\mathcal{Y}_{\theta}^{\mathrm{ice}}\text{ for }\theta=1-\nicefrac{{1}}{{p}},\text{ and }\\ \mathcal{Y}_{\beta}&=\mathrm{H}^{{2\beta},q}_{\mathrm{per},\mathrm{b.c.}}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}(G)^{4}.\end{aligned}\right.$
@@
As the spaces with horizontal periodicity embed continuously into the ones without, we omit the corresponding subscript in the remainder of the proof for convenience. For most terms, we refer to [9, Lemma 6.2], because the procedure from this lemma carries over to the present situation. Since the velocity of the atmospheric wind now acts as a variable, we verify the respective Lipschitz estimate separately. To this end, we first observe that for suitable $g_{1}$, $g_{2}$, it holds that
@@
Thanks to $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$, there is $\eta>0$ with $2-\nicefrac{{2}}{{p}}-\eta-\nicefrac{{2}}{{q}}\geq 1-\nicefrac{{2}}{{2q}}$, so [49, Theorem 4.6.1] yields that
(5.4) $\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}(G)\hookrightarrow\mathrm{B}^{2-\nicefrac{{2}}{{p}}-\eta}_{q2q}(G)\hookrightarrow\mathrm{H}^{1,2q}(G),\enspace\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}({\Omega}_{\mathrm{atm}})\hookrightarrow\mathrm{B}^{2-\nicefrac{{2}}{{p}}-\eta}_{q2q}({\Omega}_{\mathrm{atm}})\hookrightarrow\mathrm{W}^{\nicefrac{{1}}{{q}}+\nicefrac{{\eta}}{{2}},2q}({\Omega}_{\mathrm{atm}}).$
@@
Concerning $\mathrm{A}_{\omega}^{\mathrm{ice}}$, we refer to [9, Lemma 6.2] except for $C_{\mathrm{o,i}}(h)\partial_{\nu_{\Gamma_{o}}}$ which is new in comparison to the considerations in [9]. The assumption $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$ also implies continuity of $\partial_{\nu_{\Gamma_{o}}}$ from $\mathrm{H}^{{2},q}({\Omega}_{\mathrm{ocn}})$ to $\mathrm{L}^{q}(G)$. A similar argument as above for the atmospheric wind terms exhibits that

### v4

@@ 4.1. The hydrostatic Helmholtz projection and the hydrostatic Stokes operator
[…]
@@
We now introduce the associated hydrostatic Helmholtz or Leray projection $\mathcal{P}\colon\mathrm{L}^{q}(\Omega)^{2}\to{\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)$. It is a bounded, linear projection which annihilates the pressure term, i. e., $\mathcal{P}(\nabla_{\mathrm{H}}\pi)=0$, and the space $\mathrm{L}^{q}(\Omega)$ can be decomposed into $\mathrm{L}^{q}(\Omega)^{2}={\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)\oplus\{\nabla_{\mathrm{H}}\pi\colon\pi\in\widehat{\mathrm{W}}^{1,q}(G)\}$ along the hydrostatic Helmholtz projection $\mathcal{P}$. Furthermore, the Helmholtz projection commutes with the time derivative, i. e., $\partial_{t}\mathcal{P}=\mathcal{P}\partial_{t}$. It is related to the two-dimensional classical Helmholtz projection $\mathcal{P}_{\mathrm{H}}$ by
@@ 4.2. Hibler’s sea ice operator
Proceeding as in [8, Section 3], we introduce the map $\mathbb{S}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}^{2\times 2}$ with
@@
from [8, Section 3], where $i=1,2$ and $\mathrm{D}_{m}=-\mathrm{i}\partial_{m}$. The principal part of $\mathcal{A}^{\mathrm{H}}$ has coefficients given by
@@
$\mathcal{A}^{\mathrm{H}}(v_{0}){v}_{\mathrm{ice}}]_{i}=\sum_{j,k,l=1}^{2}a_{ij}^{kl}(\varepsilon_{0},h_{0},a_{0})\mathrm{D}_{k}\mathrm{D}_{l}v_{\mathrm{ice},j}+\frac{1}{2{\rho}_{\mathrm{ice}}h_{0}\triangle_{\delta}(\varepsilon_{0})}\sum_{j=1}^{2}(\partial_{j}P(h_{0},a_{0}))(\mathbb{S}\varepsilon({v}_{\mathrm{ice}}))_{ij}.$
The $\mathrm{L}^{q}$-realization $\mathrm{A}^{\mathrm{H}}(v_{0})$ of $\mathcal{A}^{\mathrm{H}}(v_{0})$ on $G=(0,1)\times(0,1)$ is then defined by $\mathrm{A}^{\mathrm{H}}(v_{0})]{v}_{\mathrm{ice}}:=[\mathcal{A}^{\mathrm{H}}(v_{0})]{v}_{\mathrm{ice}}$ for ${v}_{\mathrm{ice}}\in D(\mathrm{A}^{\mathrm{H}}(v_{0}))=\mathrm{H}_{\mathrm{per}}^{{2},q}(G)^{2}$.
Comparing this definition of Hibler’s operator to the one presented in [8, Section 3], we observe that the sign is changed and $\nicefrac{{1}}{{{\rho}_{\mathrm{ice}}h}}$ is included in the definition for the sake of consistency with the hydrostatic Stokes operator and to ease notation. We also introduce the lower-order terms
@@ 4.3. A new formulation of the coupled system
$\left\{\begin{aligned} \mathrm{A}^{\mathrm{atm}}{v}_{\mathrm{atm}}&={\mathcal{P}}_{\mathrm{atm}}\Delta{v}_{\mathrm{atm}},\\ D(\mathrm{A}^{\mathrm{atm}})&=\left\{{v}_{\mathrm{atm}}\in\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}}):\partial_{\nu_{\Gamma_{u}}}{v}_{\mathrm{atm}}=0\text{ on }{\Gamma_{u}}\text{ and }\partial_{\nu_{\Gamma_{i}}}{v}_{\mathrm{atm}}=0\text{ on }{\Gamma_{i}}\right\}\end{aligned}\right$.
@@
$\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}{v}_{\mathrm{ocn}}={\mathcal{P}}_{\mathrm{ocn}}\Delta{v}_{\mathrm{ocn}},\kern 5.0ptD(\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}})=\{{v}_{\mathrm{ocn}}\in{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}}):{\mathcal{P}}_{\mathrm{ocn}}\Delta{v}_{\mathrm{ocn}}\in{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\},$
@@
(4.3) $\left\{\begin{aligned} \mathrm{A}^{\mathrm{ocn}}_{0}{v}_{\mathrm{ocn}}&={\mathcal{P}}_{\mathrm{ocn}}\Delta{v}_{\mathrm{ocn}},\\ D(\mathrm{A}^{\mathrm{ocn}}_{0})&=\left\{{v}_{\mathrm{ocn}}\in\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}}):\mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}=0\text{ on }{\Gamma_{b}}\text{ and }\mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}=0\text{ on }{\Gamma_{o}}\right\}\end{aligned}\right$.
@@
(4.4) $\mathcal{A}(v_{\mathrm{ice},0},h_{0},a_{0})$ $=\begin{pmatrix}\mathrm{A}^{\mathrm{atm}}&0&0&0&0\\ 0&\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}&0&0&0\\ 0&-C_{\mathrm{o,i}}(h_{0})\partial_{\nu_{\Gamma_{o}}}&\mathrm{A}^{\mathrm{H}}(v_{\mathrm{ice},0},h_{0},a_{0})&\mathrm{B}_{h}(h_{0},a_{0})&\mathrm{B}_{a}(h_{0},a_{0})\\ 0&0&0&d_{\mathrm{h}}\Delta_{\mathrm{H}}&0\\ 0&0&0&0&d_{\mathrm{a}}\Delta_{\mathrm{H}}\end{pmatrix}$,
$\mathcal{X}_{1}=D(\mathcal{A})$ $=\{({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)\in D(\mathrm{A}^{\mathrm{atm}})\times D(\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}})\times D(\mathrm{A}^{\mathrm{H}})\times D(\Delta_{\mathrm{H}})\times D(\Delta_{\mathrm{H}})\colon$
@@
(4.6) $\left\{\begin{aligned} \partial_{t}u-\mathcal{A}(u)u+\mathcal{F}(u)&=f,&&t\in[0,T],\\ u(0)&=u_{0},\end{aligned}\right$.
@@ 5.1. Bounded $\mathcal{H}^{\infty}$-calculus of Hibler’s operator
The embedding below is classical, see e. g. [38, Theorem 4.6.1]. Let $p,q\in(1,\infty)$ with $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$. Then
@@
Making use of (5.1) and proceeding as in [8, Section 4], we obtain the ellipticity of $\mathcal{A}^{\mathrm{H}}(v_{0})$. For the notion of strong ellipticity and parameter ellipticity, we refer e. g. to [15, Definition 5.1].
Lemma 5.1.
Let $p,q\in(1,\infty)$ such that $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$, and let $u_{0}=(v_{\mathrm{ice},0},h_{0},a_{0})\in V^{\mathrm{ice}}$. Then the principal part of the negative Hibler operator $\mathcal{A}^{\mathrm{H}}(v_{0})$ is strongly elliptic and also parameter-elliptic of angle $\varphi_{-\mathcal{A}^{\mathrm{H}}(v_{0})}=0$.
We will now derive the bounded $\mathcal{H}^{\infty}$-calculus for the linearized $\mathrm{L}^{q}$-realization $\mathrm{A}^{\mathrm{H}}(v_{0})$.
Proposition 5.2.
Let $p,q,r,s\in(1,\infty)$ such that $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$, let $v_{0}=(v_{\mathrm{ice},0},h_{0},a_{0})\in V^{\mathrm{ice}}$, and let $\mathrm{A}^{\mathrm{H}}(v_{0})$ denote the $\mathrm{L}^{q}$-realization of the linearized Hibler operator. Then there is $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, $\mathrm{A}^{\mathrm{H}}(v_{0})+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathrm{L}^{r}(G)^{2}$ with angle $\Phi_{-\mathrm{A}^{\mathrm{H}}(v_{0})+\omega}^{\infty}<\nicefrac{{\pi}}{{2}}$.
@@
Identifying functions on the square $G=(0,1)\times(0,1)$ endowed with periodic boundary conditions with functions on the torus $\mathbb{T}$, exploiting Subsection 5.1 for the ellipticity properties of the negative Hibler operator $\mathcal{A}^{\mathrm{H}}(v_{0})$ and using (5.1) in conjunction with the fact that the coefficients of $\mathcal{A}^{\mathrm{H}}(v_{0})$ depend smoothly on $\nabla_{\mathrm{H}}v_{\mathrm{ice},0}$, $\nabla_{\mathrm{H}}h_{0}$, $\nabla_{\mathrm{H}}a_{0}$, $h_{0}$ and $a_{0}$, we conclude from [16, Theorem 7.1] that there is $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, the operator $\mathrm{A}^{\mathrm{H}}(v_{0})+\omega$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathrm{L}^{r}(\mathbb{T})^{2}=\mathrm{L}^{r}(G)^{2}$ with angle $\Phi_{-\mathrm{A}^{\mathrm{H}}(v_{0})+\omega}^{\infty}<\nicefrac{{\pi}}{{2}}$. ∎
The following result is a direct consequence of Subsection 5.1, because the latter proposition implies that $\mathrm{A}^{\mathrm{H}}(v_{0})+\omega$ has bounded imaginary powers on $\mathrm{L}^{q}(G)^{2}$, see also the relation (8.7).
Corollary 5.3.
@@ 5.2. Estimates of the nonlinearities
[…]
@@
(5.3) $\left\{\begin{aligned} \mathcal{Y}_{0}&={\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})\times{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{L}^{q}(G)^{4},\\ \mathcal{Y}_{1}&=\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})\times\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{H}_{\mathrm{per}}^{{2},q}(G)^{4},\\ \mathcal{Y}_{\theta}&=\mathrm{B}^{{2\theta}}_{{q}{p},\mathrm{per}}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})\times\mathrm{B}^{{2\theta}}_{{q}{p},\mathrm{per}}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{B}^{{2\theta}}_{{q}{p},\mathrm{per}}(G)^{4},\\ \mathcal{Y}_{\gamma}&=\mathcal{Y}_{\theta}^{\mathrm{atm}}\times\mathcal{Y}_{\theta}^{\mathrm{ocn}}\times\mathcal{Y}_{\theta}^{\mathrm{ice}}\text{ for }\theta=1-\nicefrac{{1}}{{p}},\text{ and }\\ \mathcal{Y}_{\beta}&=\mathrm{H}^{{2\beta},q}_{\mathrm{per},\mathrm{b.c.}}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{H}_{\mathrm{per}}^{{2\beta},q}(G)^{4}.\end{aligned}\right$.
@@
As the spaces with horizontal periodicity embed continuously into the ones without, we omit the corresponding subscript in the remainder of the proof for convenience. For most terms, we refer to [8, Lemma 6.2], because the procedure from this lemma carries over to the present situation. Since the velocity of the atmospheric wind now acts as a variable, we verify the respective Lipschitz estimate separately. To this end, we first observe that for suitable $g_{1}$, $g_{2}$, it holds that
@@
Thanks to $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$, there is $\eta>0$ with $2-\nicefrac{{2}}{{p}}-\eta-\nicefrac{{2}}{{q}}\geq 1-\nicefrac{{2}}{{2q}}$, so [38, Theorem 4.6.1] yields that
(5.4) $\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}(G)\hookrightarrow\mathrm{B}^{2-\nicefrac{{2}}{{p}}-\eta}_{q2q}(G)\hookrightarrow\mathrm{H}^{1,2q}(G),\kern 5.0pt\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}({\Omega}_{\mathrm{atm}})\hookrightarrow\mathrm{B}^{2-\nicefrac{{2}}{{p}}-\eta}_{q2q}({\Omega}_{\mathrm{atm}})\hookrightarrow\mathrm{W}^{\nicefrac{{1}}{{q}}+\nicefrac{{\eta}}{{2}},2q}({\Omega}_{\mathrm{atm}}).$
@@
Concerning $\mathrm{A}_{\omega}^{\mathrm{ice}}$, we refer to [8, Lemma 6.2] except for $C_{\mathrm{o,i}}(h)\partial_{\nu_{\Gamma_{o}}}$ which is new in comparison to the considerations in [8]. The assumption $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$ also implies continuity of $\partial_{\nu_{\Gamma_{o}}}$ from $\mathrm{H}^{{2},q}({\Omega}_{\mathrm{ocn}})$ to $\mathrm{L}^{q}(G)$. A similar argument as above for the atmospheric wind terms exhibits that